INFRARED SPECTROSCOPY OF SIZE-SELECTED CARBOCATIONS I.: C_{3}H_{3}^{+}, C_{3}H_{5}^{+}, AND C_{4}H_{9}^{+}

GARY E. DOUBERLY, \underline{ALLEN M. RICKS}, and MICHAEL A. DUNCAN, Department of Chemistry, University of Georgia, Athens, Georgia 30602..

Cold, rare gas tagged, gas phase carbocations are produced in a pulsed electric discharge supersonic expansion cluster source. The infrared spectra of the size-selected species are obtained via infrared photodissociation spectroscopy. The structure and spectroscopy (700-4000 cm^{−1}) of the propargyl, cyclopropenyl, allyl, 2-propenyl, and t-butyl ions will be discussed. The infrared spectral signatures of the title ions are compared to the interstellar unidentified infrared emission bands.